solve the following equation for 0«x«360

a) 6 cos²x + sin x - 4

b) 9 tan x + tan²x + 5 sec² x - 3


1
Expert's answer
2021-01-19T03:46:28-0500

a)"6\cos^2 x+\sin x-4=0"

"6(1-\sin^2 x)+\sin x-4=0"

"-6\sin^2 x+\sin x+2=0"

This quadratic equation has solutions "\sin x=-\frac{1}{2}" and "\sin x=\frac{2}{3}"

The equation "\sin x=-\frac{1}{2}" has solutions "210^\circ" and "330^\circ" in "[0^\circ,360^\circ]", and since "\frac{2}{3}>0", the equation "\sin x=\frac{2}{3}" has solutions "\sin^{-1}\frac{2}{3}" and "180^\circ-\sin^{-1}\frac{2}{3}" in "[0^\circ,360^\circ]"

b)"9\tan x+\tan^2 x+5\sec^2 x-3=0"

"9\tan x+\tan^2 x+5(1+\tan^2 x)-3=0"

"6\tan^2 x+9\tan x+2=0"

This quadratic equation has solutions "\tan x=-\frac{9-\sqrt{33}}{2}" and "\tan x=-\frac{9+\sqrt{33}}{2}"

Since "-\frac{9-\sqrt{33}}{2}<0", the equation "\tan x=-\frac{9-\sqrt{33}}{2}" has solutions "360^\circ-\tan^{-1}\frac{9-\sqrt{33}}{2}" and "180^\circ-\tan^{-1}\frac{9-\sqrt{33}}{2}" in "[0^\circ,360^\circ]"

Similarly the equation "\tan x=-\frac{9+\sqrt{33}}{2}" has solutions "360^\circ-\tan^{-1}\frac{9+\sqrt{33}}{2}" and "180^\circ-\tan^{-1}\frac{9+\sqrt{33}}{2}"

Answer: a)"210^\circ, 330^\circ, \sin^{-1}\frac{2}{3}, 180^\circ-\sin\frac{2}{3}"

b)"360^\circ-\tan^{-1}\frac{9-\sqrt{33}}{2}, 180^\circ-\tan^{-1}\frac{9-\sqrt{33}}{2},"

"360^\circ-\tan^{-1}\frac{9+\sqrt{33}}{2}, 180^\circ-\tan^{-1}\frac{9+\sqrt{33}}{2}"